3. (a) A rope is wrapped tightly around a wheel with a radius of 3 feet. If the radius of the wheel is increased by 2 feet to a radius of 5 feet, by how much must the rope be lengthened to fit around the wheel? (Round your answer to two decimal places.)
ft

Answer by rfer(16322)

6*3.14=18.84
10*3.14=31.4
31.4-18.84=12.56 longer